OpenStudy (anonymous):

Solve on the interval [0, 2pi) 1- cos theta = 2- sqt3/2

OpenStudy (anonymous):

A. 2pi/3, 4pi/3 B. pi/6, 5pi/6 C. pi/6, 11pi/6 D. 7pi/6 , 11pi/6

OpenStudy (anonymous):

say cos theta is x 1-x = 2-3^0.5/2......is this the qn?

OpenStudy (anonymous):

i think yes

OpenStudy (anonymous):

C

OpenStudy (anonymous):

sry......x=-1+cos(30)....